OpenStudy (anonymous):

WILL GIVE FAN AND MEDAL! Tell whether the given number is a solution of each equation. 74. 9 = 2a + 3; 3 76. 4.4r – 2.8 = 1.6; 1

OpenStudy (anonymous):

Please help! @mathmale

OpenStudy (anonymous):

they are 2 different equations

OpenStudy (mathmale):

Next time, please type in only one problem at a time. Do you mean 74. 9 = 2a + 3; 3?

OpenStudy (anonymous):

Yes. :)

OpenStudy (acxbox22):

for each question plug in the given value in place of the variable and see if both sides are equal and yeah post each question separately...

OpenStudy (mathmale):

Then please substitute the value 3 for the letter a. Is the resulting equation true or false?

OpenStudy (mathmale):

Is 9 = 2(3) + 3 true or false?

OpenStudy (anonymous):

both are correct

OpenStudy (mathmale):

OK. Is 3 a solution of the first equation?

OpenStudy (anonymous):

What? 9 is

OpenStudy (mathmale):

No. think what "solution" means. You are substituting the value 3 for the variable 'a' in the first equation. When you do that, you find that the equation is true. We say that 3 is a ... ? of the equation.

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Please provide the missing word. The first equation is true if we substitute a = 3. We therefore call a = 3 a " ....................... " of the equation.

OpenStudy (anonymous):

variable

OpenStudy (mathmale):

a is a variable, yes. But what do the instructions for this problem as you to do? a = 3 is a particular value of the variable 'a'. since the equation is true when a = 3, we call that a = 3 a " ................ " of the equation. Go back and read the problem statement, please.

OpenStudy (anonymous):

given number

OpenStudy (anonymous):

solution

OpenStudy (mathmale):

Yes. Take my word for this: you have to understand that word "solution" as you will see it again and again in math.

OpenStudy (mathmale):

r = 1 is a " ............... " of ..... ?

OpenStudy (mathmale):

What happens when you substitute r = 1 into the second equation?

OpenStudy (anonymous):

whatttttttttttttttt

OpenStudy (mathmale):

Same thing as before. If you substitute r=1 into the second equation, the equation is true. That's what you yourself told me. Therefore, we call r = 1 a ' ................ ' of ... what ... ?

OpenStudy (anonymous):

i give up

OpenStudy (mathmale):

Review: "Please provide the missing word. The first equation is true if we substitute a = 3. We therefore call a = 3 a " ....................... " of the equation. a is a variable, yes. But what do the instructions for this problem as you to do? a = 3 is a particular value of the variable 'a'. since the equation is true when a = 3, we call that a = 3 a " ................ " of the equation. Go back and read the problem statement, please."

OpenStudy (mathmale):

Please substitute r =1 into 4.4r – 2.8 = 1.6. (You have done this before.) Is the equation now true or false?

OpenStudy (mathmale):

When you respond, I will also. Since you're not responding at the moment, I'm leaving this problem temporarily to look at problems posted by others.